OpenStudy (anonymous):

the distance travelled by a body falling freely from in the 1st, 2nd and 3rd seconds, are in the ratio a)1:2:3 b)1:3:5 c)1:4:9 d)3:2:1

OpenStudy (luffingsails):

adhira, A free-falling body will accelerate. This should help you answer the question. Only one of the answers indicates this behavior.

OpenStudy (anonymous):

What do you mean by 1st, 2nd and 3rd second? Can you explain your question a little more.

OpenStudy (anonymous):

pls can u explain that i haven't the answer key of this question upsilon

OpenStudy (anonymous):

but there u=0

OpenStudy (luffingsails):

given gravity =32 ft/sec/sec d = 16 t^2 So, distances for seconds 1, 2, and 3 = 16, 64, and 144 respectively. The ratio that represents these is 1:4:9.

OpenStudy (anonymous):

yeah i have the same ans 1:4:9 it's my exam question today

OpenStudy (anonymous):

\[v = u +gt \] Now at \(t = 0 \), \(u=0\). \[v = gt \implies v = 10\] \[s = ut + \frac{1}{2}*g*t^2\] \[s _1= 5 m \] For the second time interval the final velocity of 1st interval will be initial velocity \[s_2 = 10 + 5\] \[s _2= 15\] Now final velocity for second interval \(v = 10 + gt\), \(v=20\) \[s_3 = 20 + 5\] \[s_1:s_2:s_3 = 1:3:5\]

OpenStudy (anonymous):

thanks thanks thanks thanks UPSILON U DONE A VERY VERY GOOD JOB

OpenStudy (anonymous):

Naruto San ...Sensei here

OpenStudy (anonymous):

Yea I am with upsilon